The oxygen requirement in an aeration basin is 1500 lb/hr. The volume of the basin is 3 MG. Temperature is 30 degrees Celsius. The compressor efficiency is 80%. The elevation is 2500 ft and the diffusors are submerged to a depth of 12 ft. Find the power required by the compressor.
Answer #1

mass flow rate = 1500 lb/hr = 0.000125998 * 1500 kg/s

= 0.1889968 kg/s

total height = 2500 + 12 = 2512 ft = 0.3048 * 2512 = 765.6576 m

Power required by compressor = (mass flow rate * height * g)/ (efficiency of compressor)

= (0.1889968 * 765.6576 * 9.81)/.8

=1774.46 W

= 1.77446 KW Answer